OpenStudy (anonymous):

Which one of the following points satisfies the equation 2x+3y=8 A) (-1,3) B) (1,4) C) (-2,4)

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

just substitute x & y in equation: A) -2+9=7 not 8 doesn't satisfies B) 2+12=14 not 8 C) -4+12= 8 = 8 True! this is your point!